OpenStudy (anonymous):

Hey what is the gcf of 38 and 51 @skateboarding

OpenStudy (anonymous):

none

OpenStudy (anonymous):

38=19*2;51=17*3

OpenStudy (anonymous):

lovemath's factorizations are correct. You would say that 1 is the gcf.